# AdaFold: Adapting Folding Trajectories of Cloths via Feedback-loop Manipulation

### v1

@@ Abstract
We present AdaFold, a model-based feedback-loop framework for optimizing folding trajectories. AdaFold extracts a particle-based representation of cloth from RGB-D images and feeds back the representation to a model predictive control to re-plan folding trajectory at every time-step. A key component of AdaFold that enables feedback-loop manipulation is the use of semantic descriptors extracted from visual-language models. These descriptors enhance the particle representation of the cloth to distinguish between ambiguous point clouds of differently folded cloths. Our experiments demonstrate AdaFold’s ability to adapt folding trajectories to cloths with varying physical properties and generalize from simulated training to real-world execution.
@@ I Introduction
Generalizing robotic manipulation skills requires adapting to object variations such as pose, shape, and physical properties [1]. Feedback-loop manipulation represents a class of methods to adapt to these variations [2]. The effectiveness of this class of methods, however, is heavily contingent upon the robot’s ability to accurately perceive and track the state of the object throughout the manipulation. Within the realm of deformable objects, such as cloth, feedback-loop manipulation remains under-explored due to challenges in state estimation and dynamics modelling [3].
Recent advances in learning cloth dynamics have led to model-based methods to predict and plan pick-and-place interactions for folding and flattening [4, 5]. Despite the progress, these methods rely on open-loop planning and pre-defined manipulation trajectories due to the practical challenges of estimating and tracking cloth states during manipulation. A promising approach to mitigate these challenges is integrating semantic knowledge, obtained from pre-trained visual-language models (VLMs), into potentially ambiguous, but easy to track, representations such as point clouds [6]. Nevertheless, the effectiveness of VLMs in accurately interpreting a wide range of cloth configurations is yet to be fully explored.
@@
Figure 1: Overview of AdaFold for feedback-loop manipulation of cloths. Given a set of pick-and-place positions $(x_{pick},x_{place})$, AdaFold optimizes the best folding action $a^{*}_{t}$ at each time-step $t$. RGB-D observations from different calibrated cameras are used to extract point cloud representations with semantic descriptors. The semantic descriptors Upper and Bottom are provided by VLMs and refined by our ensemble mask selection process. The optimal folding action $a^{*}_{t}$ is obtained with MPC, which uses the forward and adaptation modules $f_{\theta}$ and $g_{\psi}$ to evaluate the candidate trajectories $a^{n}$ (light blue) and update the optimal control sequence $a^{*}$ (dark blue).
We propose AdaFold, a model-based framework for feedback-loop manipulation of cloth to optimize folding trajectories. AdaFold relies on particle-based state representation and a learned model of the cloth to optimize the best sequence of folding actions with model-predictive control (MPC) [7]. To adapt to cloth variations, we perform the manipulation in a feedback-loop fashion by re-planning the folding trajectory after every time-step. We further propose to use pre-trained VLM to extract semantic descriptors of the upper and the bottom layers of the cloth from RGB images. To overcome the limitations of VLMs in interpreting semantic and spatial relationships inherent in deformable objects [8], we introduce a mask selection step based on an ensemble of prompts to accurately extract the semantic descriptors of the points.
@@
  • A method to embed semantic descriptors into a point cloud representation of cloth by leveraging pre-trained VLMs.
@@ II Related Work
II-A Cloth Manipulation
Model-free learning is a class of methods that circumvents the challenges in state estimation and dynamics modeling. This technique has been successfully used to tackle several cloth manipulation tasks such as folding or flattening, by finding the best sequence of pick-and-place positions [10, 11, 12, 13, 14]. More recently, [15] proposed a model-free visual feedback policy to fold cloths in half, successfully adapting the manipulation trajectory to three real-world cloths. To address the sample inefficiency of model-free learning [16], the authors relied on human expert demonstrations. Model-based techniques, on the other hand, have shown promise in sample-efficient learning [17, 4, 5]. These approaches often adopt a top-down camera view, which effectively minimizes the self-occlusions of the cloth but restricts the visibility of the cloth’s intermediate states during manipulation. This limits the applicability of feedback-loop strategies. In this work, we leverage different camera views to reduce occlusions and propose a model-based approach for feedback-loop manipulation.
II-B State Representation
@@
While image-based representations alleviate the need for explicit state estimation and tracking of the cloth [18, 19], they are sensitive to variations in color, brightness, and camera perspective. Consequently, a large amount of data is required to train a policy robust to these variations [20]. Particle-based representations, such as graphs, have shown the potential to decrease reliance on color and visual randomization due to their invariance to visual features, demonstrating better generalization to novel cloth shapes and textures [5, 4]. However, tracking graph representations of deformable objects is still an open problem [21].
@@
Point clouds, on the other hand, demand less computational effort for perception tasks due to their unstructured nature and have shown success in assistive dressing tasks [13]. Nevertheless, in situations where clothes significantly self-occlude, point cloud representations become ambiguous as different layers of the cloth cannot be distinguished based solely on the observable set of points. While classical computer vision approaches such as a Harris Corner Detector [22] or a wrinkle-detector [23] can be used for detecting cloth features, they are typically not robust to variations of texture, lightning conditions, and non-static observations. This study tackles these perception challenges by integrating semantic descriptors, derived from RGB observations through pre-trained VLMs, with point cloud representations. While VLMs have shown promise in recognizing rigid objects and clothing to some extent [24], their ability to interpret diverse cloth information remains largely unexplored.
@@ III Problem Formulation
where $T$ is the control horizon and $\mathcal{J}(\tau_{0:T})$ denotes the objective value of the trajectory $\tau_{0:T}$ The control input is updated at every time step in a feedback-loop fashion using MPC with horizon $H$. Commonly, the cloth dynamics $f$ and the physical properties $\xi$ are not precisely known and vary across different objects [27]. Thus, similarly to [28, 29], we jointly learn 1) an adaptation module $g_{\psi}$ to encode a recent history of observations into a latent representation $z_{t}$ of the physical properties and adapt to different objects, 2) an approximate model $f_{\theta}$ of the cloth dynamics conditioned on the latent representation. These models are respectively parameterized by $\theta$ and $\psi$. Our framework couples cloth perception with data-driven trajectory optimization for Adaptive Folding, which we denote AdaFold.
@@ IV Cloth Perception
The cloth perception consists of processing RGB-D observations into semantically labelled point cloud $P=P^{U}\cup P^{B}$ representing the cloth.
IV-A Labeled Point Cloud
Given an RGB-D image $I_{t}$, we first find the segmentation mask of the cloth from the RGB observation using the promptable and fine-grained segmentation modules Grounding-DINO [6] and Segment Anything (SAM) [30]. Then, we transform the masked depth into a point cloud corresponding to the observable points of the cloth $P_{t}$ using the camera’s intrinsic matrix. To assign the semantic labels to $P_{t}$, we further segment the RGB image into upper and bottom layers. We repeat this process for images $\{I_{t}^{1},I_{t}^{2}\}$ recorded from two calibrated cameras to reduce the effect of self-occlusions. The resulting point clouds are then transformed into a common reference frame, for example, the robot base frame. This step aggregates the point clouds to derive the final cloth state $P_{t}=P_{t}^{U}\cup P_{t}^{B}$, which can be further voxelized to make the density of the points uniform across the cloth surface. An example segmentation of the different parts of the cloth and the point cloud reconstruction is visualized in Fig. 1.
Figure 2: Example of the mask selection process. Given all the masks obtained from Grounded SAM with the ensemble of prompts ${C=\{C_{1},C_{2},C_{3},C_{4}\}}$, our selection process correctly filters the mask to obtain the corresponding masks of the full, bottom and upper layers of the cloth.
IV-B Ensemble Mask Selection
We initially relied on SAM’s confidence values to select the masks for the upper and bottom layers of the cloth. However, we were unable to identify single prompts that consistently yielded correct masks throughout every step of the manipulation, despite trying a diverse set of prompts (e.g. “bottom layer of the cloth”, “bottom black layer of the cloth”, “bottom half of the rectangular cloth” amongst others). Common failure instances involved improper segmentation of either the upper or lower half of the cloth, resulting in a complete cloth mask rather than the intended half, or conversely, segmenting only one half when the entire cloth was desired. To address this challenge, we introduce an alternative mask selection strategy that considers an ensemble of prompts ${C=\{C_{1},\ldots,C_{n}\}}$, with $n$ the number of prompts, and aggregates the obtained masks based on the following heuristics:
Full: We obtain the mask of the full cloth $M_{f}$ by aggregating all the masks from the set of prompts $C$, where the aggregation corresponds to a logical OR operation.
Bottom: We assume that the bottom layer has a predefined color. Leveraging this assumption, we select the masks corresponding to the bottom layer by initially imposing a color threshold $r$ on each resulting masked RGB image. Subsequently, we aggregate only the masks with many pixels satisfying this threshold, thereby obtaining ${M_{b}}$. Differently from standard color threshold techniques, we utilize this threshold as a soft voting mechanism without the need to accurately choose $r$.
Upper: Given the two previous steps, the upper mask is easily obtained as the difference between the full and the bottom mask: $M_{u}=M_{f}-M_{b}$.
In Fig. 2, we present an example of the masks selected by our procedure.
@@ V-A Learned Model
where $\hat{P}_{t+m}=f_{\theta}(\hat{P}_{t+m-1},a_{t},z_{t})$, $z_{t}=g_{\psi}(\tau_{t-K:t-1})$ and $M$ represents the number of steps in the future to be predicted. We use the MSE over the Chamfer loss as we empirically observed more stable training without compromising accuracy. Finally, we assume that ${\hat{P}^{B}_{t+1}={P}^{B}_{t}}$ due to the friction between the cloth and the table, preventing the bottom half from slipping during manipulation. Thus, it is sufficient for the model to predict only $\hat{P}^{U}_{t+1}$.
@@ 15       $\text{Execute}\quad a^{*}_{t}$
Cost function: The cost function is designed to select actions that maximize the alignment between the two halves of the cloth. This alignment is quantitatively assessed using the Intersection over Union (IoU) metric, calculated based on the areas occupied by each half $P^{U}$ and $P^{B}$ when projected to the folding plane. We additionally incorporate in the cost function a term discouraging actions resulting in large displacements of the cloth. Specifically, the cost function is defined as a weighted sum of two cost terms:
@@
Constrained random sampling: To reduce the action search space, we constrain the random candidate control sequence sampling at time $t$ by sampling $T-t$ action directions $\mathbf{d}$ similar to the pick-and-place direction $\mathbf{d}_{pp}$. Specifically, we impose the cosine similarity (CS) to satisfy $\text{CS}(\mathbf{d},\mathbf{d}_{pp})\geq 0$. The actions are then defined as $a=(\mathbf{d}/\lVert\mathbf{d}\lVert)v$ where $v$ is a fixed value specifying the action norm.
Place reaching: If the termination of the sampled trajectory is not sufficiently close to the place position, we use a heuristic to extend the path. This heuristic continues the trajectory along a linear path towards $x_{\text{place}}$. Fig. 1 shows an example of candidate trajectories (light blue) terminating in the place position. While extending the trajectory to the place position is a reasonable assumption since $x_{\text{place}}$ is known, it could be removed by incorporating an attractor towards the place position into the cost function, with the disadvantage of an additional hyperparameter to tune.
@@ VI-A Half folding task
Overview: We consider the half folding task proposed in  [9]. In this task, the agent controls a square or rectangular cloth placed approximately in the centre of the environment at the beginning of each episode. We assume the agent starts from a grasped state, while the objective is to fold the cloth in half by aligning the corners.
Success Criteria: We evaluate the success of the folding execution by computing the 2D intersection over union (IoU) between the two halves of the cloth as described in Section V-B. We use it as a reference for the bottom half of the points observed at time $t=0$ as throughout the manipulation, the bottom half will be occluded.
@@ VI-B Simulation
2 $17\times 25$ Polyester Woven-Plain 0.56 0.4
@@
Figure 3: Visualization of the real-world set-up: two Realsense D435 cameras capturing different views of the scene, and the dataset composed of 6 rectangular cloths.
@@ VI-C Real-world
Perception: We collected observations from two calibrated cameras. The prompt used for grounded SAM to obtain the semantic masks were: [[[[ “black bottom half cloth”, “upper colored half cloth”, “grasped central cloth”, “rectangular cloth” ]]]]. We further filtered the point cloud to remove outliers and voxelize it with a voxel size of $0.8$cm.
Dataset: We utilized six rectangular cloths with different material, construction and physical properties [35] (see Fig. 3), none of them matching the one used in the simulation. Specifically, five of the six cloths shared the same shape but differed in physical properties, whereas the sixth cloth was slightly larger than the rest. We report the cloth properties in Table I. We used a black cover throughout the evaluation to facilitate the distinction between the upper and bottom layers of the cloth.
@@ VII-A Baselines
We compare AdaFold against four different baselines. The first baseline is a fixed triangular trajectory (Triangular) selected as the best performing in simulation when compared to a linear trajectory. The second is a random baseline (Random), which, at each time step, randomly selects an action among all the candidate actions. The third baseline is an open-loop version of AdaFold (AdaFold-OL) that optimizes the trajectory once at the beginning of the folding. The final baseline we compared to was the model-free learning method DDPG [36]. For a fair comparison, we trained DDPG offline on the same dataset as AdaFold to learn a state-action value function. At test time, we used the learned value function to select the best action among the candidates (DDPG-Critic). The reward function and the model architecture for this baseline correspond to the one chosen for our approach, where the critic architecture and hyperparameters match the ones of AdaFold. As the output of the critic has the same dimension as the number of input points, to extract the value from the output of the segmentation branch, we select the first point of the network output as similarly done in [13].

### v3

@@ Abstract
We present AdaFold, a model-based feedback-loop framework for optimizing folding trajectories. AdaFold extracts a particle-based representation of cloth from RGB-D images and feeds back the representation to a model predictive control to re-plan folding trajectory at every time-step. A key component of AdaFold that enables feedback-loop manipulation is the use of semantic descriptors extracted from geometric features. These descriptors enhance the particle representation of the cloth to distinguish between ambiguous point clouds of differently folded cloths. Our experiments demonstrate AdaFold’s ability to adapt folding trajectories of cloths with varying physical properties and generalize from simulated training to real-world execution. Videos can be found on our project website: https://adafold.github.io
Index Terms:
Manipulation Planning; Perception for Grasping and Manipulation; RGB-D Perception; Semantic Scene Understanding
@@ I Introduction
Generalizing robotic manipulation skills requires adapting to object variations such as pose, shape, and physical properties [1]. Feedback-loop manipulation represents a class of methods to adapt to these variations. The effectiveness of this class of methods, however, is heavily contingent upon the robot’s ability to accurately perceive and track the state of the object throughout the manipulation. Within the realm of deformable objects, such as cloth, feedback-loop manipulation remains under-explored due to challenges in state estimation and dynamics modelling [2, 3].
Recent advances in learning cloth dynamics have led to model-based methods to plan pick-and-place interactions for folding and flattening [4, 5]. Despite the progress, these methods rely on open-loop planning and pre-defined manipulation trajectories due to the practical challenges of estimating and tracking cloth states during manipulation. A promising approach to mitigate these challenges is integrating semantic knowledge based on geometric features [6] into potentially ambiguous, but easy to track, representations such as point clouds.
@@
Figure 2: Overview of AdaFold for feedback-loop manipulation of cloths. Given a set of pick-and-place positions $(x_{pick},x_{place})$, AdaFold optimizes the best folding action $a^{*}_{t}$ at each time-step $t$. RGB-D observations from different calibrated cameras are used to extract point cloud representations with semantic descriptors. The semantic descriptors Upper and Bottom are obtained based on geometric features following [7]. The optimal folding action $a^{*}_{t}$ is obtained with MPC, which uses the forward and adaptation modules $f_{\theta}$ and $g_{\psi}$ to evaluate the candidate trajectories $a^{n}$ (light blue) and update the optimal control sequence $a^{*}$ (dark blue).
We propose AdaFold, a model-based framework for feedback-loop manipulation of cloth to optimize folding trajectories. AdaFold relies on particle-based state representation and a learned model of the cloth to optimize the best sequence of folding actions with model-predictive control (MPC) [8]. To adapt to cloth variations, we perform the manipulation in a feedback-loop fashion by re-planning the folding trajectory after every time-step. We further propose to use parameterized shape models for cloths [7] to extract semantic descriptors of the upper and the bottom layers of the cloth from RGB images. To overcome the challenge of tracking the upper and bottom layers through manipulation, we leverage recent advances in video tracking [9] to track the respective masks.
@@
  • A method to embed semantic descriptors into a point cloud representation of cloth by leveraging geometric features along pre-trained video tracking techniques.
@@ II Related Work
II-A State Representation
@@
Geometric features, such as corners or landmark points, are effective representations when the cloth lies flat on a surface, but fail for more crumpled cloth configurations [11]. While image-based representations alleviate the need for explicit state estimation and tracking [12, 13], they are sensitive to variations in color, brightness, and camera perspective. Particle-based representations, such as graphs, are less sensitive to visual clues and achieve better generalization to novel cloth shapes and textures [5, 4]. Yet, tracking graph representations is still a challenge [14]. Hierarchical and bottom-up representations [15] have not yet been demonstrated in the context of cloth manipulation.
@@
Point clouds, on the other hand, demand less computational effort for perception tasks due to their unstructured nature and have shown success in assistive dressing tasks [16]. Nevertheless, in situations with significant self-occlusions, point cloud representations become ambiguous as different layers of the cloth cannot be distinguished based solely on the observable set of points. While classical computer vision approaches such as a Harris Corner Detector [17] or a wrinkle-detector [18] can be used for detecting cloth features, they are typically not robust to variations of texture, lightning conditions, and non-static observations. This study tackles these perception challenges by augmenting point cloud representations with semantic descriptors, which we derive from RGB observations through geometric features.
II-B Cloth manipulation
Heuristic-based methods for cloth manipulation are viable approaches for cloth folding and flattening [6], although they may struggle to generalize across diverse real-world scenarios. Learning-based methods emerge as an alternative when heuristics fall short, and can be divided into two categories: model-free and model-based. Model-free learning is a class of methods that circumvents the challenges in state estimation and dynamics modeling while also scaling to the real world. This technique has been successfully used to tackle several cloth manipulation tasks, such as folding or flattening, by finding the best sequence of pick-and-place positions [19, 20, 21, 16, 22]. More recently, [23] proposed a model-free visual feedback policy to fold cloths in half, successfully adapting the manipulation trajectory to three real-world cloths. To address the sample inefficiency of model-free learning [24], the authors relied on human expert demonstrations. Model-based techniques, on the other hand, have shown promise in sample-efficient learning [25, 4, 5]. These approaches typically employ a top-down camera view to minimize self-occlusions of the cloth, which, however, restricts visibility of the cloth’s intermediate states during manipulation. Consequently, this limitation hinders the use of feedback-loop strategies. In this work, we address this limitation by leveraging side camera views to reduce occlusions and propose a model-based approach specifically designed for feedback-loop manipulation.
Figure 3: Cloth labeling: Given the cloth mask at time $t=0$, we obtain fold lines based on geometric features following [7]. These fold lines define the upper and bottom layers based on the pick and place locations.
@@ III Problem Formulation
where $T$ is the control horizon and $\mathcal{J}(\tau_{0:T})$ denotes the objective value of the trajectory $\tau_{0:T}$ The control input is updated at every time step in a feedback-loop fashion using MPC with horizon $H$. Commonly, the cloth dynamics $f$ and the physical properties $\xi$ are not precisely known and vary across different objects [28]. Thus, similarly to [29], we jointly learn 1) an adaptation module $g_{\psi}$ to encode a recent history of observations into a latent representation $z_{t}$ of the physical properties and adapt to different objects, 2) an approximate model $f_{\theta}$ of the cloth dynamics conditioned on the latent representation. These models are respectively parameterized by $\theta$ and $\psi$. Our framework couples cloth perception with data-driven trajectory optimization for Adaptive Folding, which we denote AdaFold.
@@ IV Cloth Perception
The cloth perception consists of processing RGB-D observations into a semantically labelled point cloud $P=P^{U}\cup P^{B}$ representing the cloth. Specifically, our method involves two main steps: 1) extracting at time $t=0$ the masks representing the upper and bottom layers of the cloth using its geometric features, and 2) tracking these masks with pre-trained video tracking models.
Given an RGB-D image $I_{0}$ at time step $t=0$, we start by finding the segmentation mask of the full cloth $M_{f}$ from the RGB observation using an object segmentation module. In this work, we use Grounding-DINO [30] and Segment Anything (SAM) [31] along with the prompt “cloth”. Given the binary mask of the cloth, we identify geometric features by fitting a parameterized polygon to the contour of the cloth as suggested in  [7]. These geometric features, represented by the cloth corners for the specific case of rectangular cloths, allow the definition of task-specific folds by specifying a directed line segment in the plane that partitions the polygon into two parts, one to be folded over another [6]. To specifically implement half-folding, we project the pick-and-place positions $x_{\text{pick}},x_{\text{place}}$ onto the image. These positions guide the selection of the fold segments that separate the cloth mask into upper $M_{u}$ and bottom $M_{b}$ halves, corresponding to the upper and bottom layers of the fold. Specifically, the upper mask includes $x_{\text{pick}}$ but not $x_{\text{place}}$, while the bottom mask includes $x_{\text{place}}$ but not $x_{\text{pick}}$. Following this segmentation, we transform the masked depth observations into a point cloud $P_{t}$ using the camera’s intrinsic matrix and label each point that belongs to the upper $P^{U}$and bottom $P^{B}$ layers according to the labels of the mask. An overview of this labeling process is detailed in Fig. 3.
While extracting these features is straightforward from flattened configurations, the process becomes more challenging when the cloth is manipulated, as its shape undergoes deformation. We address this problem by tracking the full $M_{f}$ and upper $M_{u}$ masks of the cloth using video tracker XMEM [9] at each time-step of the manipulation. Given these two masks, the bottom mask is easily obtained as ${M_{b}=M_{f}-M_{u}}$.
@@ V-A Learned Model
where $\hat{P}_{t+m}=f_{\theta}(\hat{P}_{t+m-1},a_{t},z_{t})$, $z_{t}=g_{\psi}(\tau_{t-K:t-1})$ , $K$ is the length of the recent history use to encode the latent representation $z_{t}$ and $M$ represents the number of steps in the future to be predicted. We use the MSE over the Chamfer loss as we empirically observed more stable training without compromising accuracy. Finally, we assume that ${\hat{P}^{B}_{t+1}={P}^{B}_{t}}$ due to the friction between the cloth and the table, preventing the bottom half from slipping during manipulation. Thus, it is sufficient for the model to predict only $\hat{P}^{U}_{t+1}$.
@@ 17      $\text{Execute}\quad a^{*}_{t}$
Cost function: The cost function is designed to select actions that maximize the alignment between the two halves of the cloth. This alignment is quantitatively assessed using the Intersection over Union (IoU) metric, calculated based on the ratio of the areas occupied by each half $P^{U}$ and $P^{B}$ when projected to the folding plane. We additionally incorporate in the cost function a term discouraging actions resulting in large displacements of the cloth. Specifically, the cost function is defined as a weighted sum of two cost terms:
@@
Constrained random sampling: To reduce the action search space, we constrain the sampling of the random candidate control sequence at time $t$ by sampling directions $\mathbf{d}$ similar to the pick-and-place direction $\mathbf{d}_{pp}$. Specifically, we impose the cosine similarity (CS) to satisfy $\text{CS}(\mathbf{d},\mathbf{d}_{pp})\geq 0$, which enforces the actions to always move towards the place location. The actions are then defined as $a=(\mathbf{d}/\lVert\mathbf{d}\lVert)v$ where $v$ is a fixed value specifying the action norm.
Place reaching: If the termination of the sampled trajectory is not sufficiently close to the place position, we use a heuristic to extend the path. This heuristic continues the trajectory along a linear path towards $x_{\text{place}}$ Fig. 2 shows an example of candidate trajectories (light blue) terminating in the place position. While extending the trajectory to the place position is a reasonable assumption since $x_{\text{place}}$ is known, it could be removed by incorporating an attractor towards the place position into the cost function, with the disadvantage of an additional hyperparameter to tune.
@@ VI-A Half folding task
Overview: We consider the half folding task proposed in  [10]. In this task, the agent controls a square or rectangular cloth placed approximately in the center of the environment at the beginning of each episode. We assume the agent starts from a grasped state, while the objective is to fold the cloth in half by aligning the corners.
Success Criteria: We evaluate the success of the folding execution by computing the 2D intersection over union (IoU) between the two halves of the cloth as described in Section V-B. As the bottom half of the cloth will be occluded in later stages of the manipulation, we use as a reference the points observed at time $t=0$.
@@ VI-B Simulation
2 $17\times 25$ Polyester Woven-Plain 0.56 0.40
@@
7 $28\times 20$ Cotton Woven-Plain 0.87 0.22
Figure 4: Visualization of the real-world set-up: two Realsense D435 cameras capturing different views of the scene, and the dataset composed of $7$ cloths.
@@ VI-C Real-world
Dataset: We utilized six rectangular cloths with different material, construction, and physical properties [36], none of them matching the one used in the simulation, and one pair of pants (see Fig. 4). Specifically, cloths $2-6$ share the same shape but differ in physical properties. Cloth $1$ retains the rectangular shape, but it is bigger than the others. Cloth $7$ differs in shape and size. We report the cloth properties in Table I.
Perception: We collected observations from two calibrated cameras. To find the geometric features from the mask of the full cloth, we fit a polygon with respectively $4$ and $7$ vertices to the rectangular and pants objects. We then constructed the point cloud based on the camera intrinsics and extrinsic parameters. We further filtered the point cloud to remove outliers and voxelize it with a voxel size of $0.8$cm. To track the masks, we used an open-source pre-trained video tracker available at  [9].
@@ VII-A Baselines
We compare AdaFold against five different baselines. The first baseline, representing most of open loop methods, is a fixed triangular trajectory (Triangular), which was selected as the best performing in simulation when compared to a linear trajectory. The second is a random baseline (Random), which, at each time step, randomly selects an action among all the candidate actions. The third and fourth baselines are ablations of AdaFold, one considering an open-loop version of AdaFold (AdaFold-OL) that optimizes the trajectory once at the beginning of the folding, and one that uses a point cloud without labels (AdaFold-NL). The final baseline we compared to was the model-free learning method DDPG [37]. For a fair comparison, we trained DDPG offline on the same dataset as AdaFold to learn a state-action value function. At test time, we used the learned value function to select the best action among the candidates (DDPG-Critic). The reward function and the model architecture for this baseline correspond to the one chosen for our approach, where the critic architecture and hyperparameters match the ones of AdaFold. As the output of the critic has the same dimension as the number of input points, to extract the value from the output of the segmentation branch, we select the first point of the network output as similarly done in [16].